Someone please help me on these questions. Thank you
There are only 5 tries for this problem ) A) The chemical energy stored in food is an example of source energy B) Source energy can be purely C) Macroscopic processes that change source encrgy always increase thermal energy as well D) Mechanical energy can never be converted back into source energy energy 19) You push a 30 kg crate, accelerating from rest to a speed of 3 ms. You use up 3000 J of your energy. How energy efficient is this process? Answer:
Answer #1

18) AC

19) work done = change in kinetic energy = 0.5 x 30 x (32) = 135J

efficiency = 135 x 100/3000 = 4.5%

20) thermal energy = 3000-135 = 2865 J